# Halo Mass Functions at High Redshift

### v1

@@ 2.4.1 Press-Schechter Theory
where $\delta_{\text{c}}\approx 1.686$ is the critical overdensity required for a region to spherically collapse into a dark matter halo. A limitation of the PS fitting function is its tendency to overestimate the number of lower-mass halos and underestimate the number of higher-mass halos compared to N-body simulations (Lacey & Cole (1994), Sheth & Tormen (1999)). Since this seminal paper many other fitting functions have been developed that aim to correct for this discrepancy as well as attempting to add additional sophistication to the modelling. In this work, we explore four other halo mass fitting functions in addition to PS.
@@ 2.4.2 Sheth, Mo & Tormen and Beyond
where $A=0.3222$, $a=0.707$ and $p=0.3$. We also test fitting functions developed by Reed et al. (2007) (Reed07) and Watson et al. (2013) (WatsonFoF refers to a fit using the Friends-of-Friends halo finder, WatsonSO refers to a fit using the Spherical Overdensity halo finder). The PS and SMT fits are widely-used and based on an analytical formalism, with no restriction on redshift or halo mass range. The Reed07, WatsonFoF and WatsonSO fits are based on simulation results, all with a redshift range of $z\,\in(0,30)$ and no explicit restriction on halo mass. While many other halo mass fitting functions exist we chose these fits as they were all calibrated across the redshift range of interest to us (i.e. $z>10$). Other fitting functions are typically calibrated for lower redshifts.
@@ 3 Results
As previously stated the goal of this paper is to compare halo number densities derived from simulation results with those derived from popular fits and to test how well the fits can predict JWST halo mass abundances compared to simulations (e.g. Boylan-Kolchin, 2023).
Figure 3: The ratios of the fitted halo number densities to the numerical halo number densities, derived from Enzo simulations and using the FOF halo finder. Differences between Enzo and the (semi-)analytical HMFs are typically less than a factor of two at $z=10.0$. Increasing only marginally out to $z=15.0$.
Figure 4: The ratios of the fitted halo number densities to the numerical halo number densities, derived from SWIFT simulations and using the FOF halo finder. Similar to the Enzo result, SWIFT shows excellent agreement with the (semi-) analytic HMFs, all fits except PS agreeing within a factor of 2 for all but the highest masses at $z=10.0$. There is a greater discrepancy present at $z=15.0$.
In Figure 1, we compare the HMFs from Enzo and SWIFT data (dashed lines) with the chosen fits (solid lines) for $z=20.0$, $z=15.0$ and $z=10.0$ using the FOF halo finder. The black rectangles indicate the high mass range depicted in more detail in Figure 2. The range of halo masses selected in the black rectangles is bounded by the approximate lowest halo masses observable by JWST at these epochs up to the most massive halo masses accessible by the numerical simulations at that redshift. Given that JWST is unlikely to be able to access galaxies at $z\sim 20$ (except perhaps via extreme lensing) we do not add zoom in rectangles for our $z=20.0$ outputs. Overlaid onto Figure 1 are yellow rectangles which act as a visual aid, depicting an estimate for the range of halo masses that host recently-observed JWST galaxies (Chakraborty et al., 2024). For a given photometric redshift $z$ and stellar mass $M_{*}$, we estimate the host halo mass in the usual way:
@@
where we set $\epsilon_{*}=0.1$. $\epsilon_{*}$ is the star formation efficiency parameter.
@@
Over the full mass range, for both simulation suites and halo finders, we find the fits considered agree well across the entire spectrum. There is, over several dexes in mass and over the redshift range between $z=10.0$ and $z=20.0$, excellent agreement between all of the HMFs. The largest discrepancies arise at the highest halo masses and the highest redshifts. Aside from the PS fit, the other fits agree closely with numerical simulations - particularly in the JWST range at $z=10.0$, where they agree within a factor of 2. Additionally, the fits improve as redshift decreases - not just between the numerical results and the (semi-)analytic results but even among the (semi-)analytic fits themselves.
As discussed, in Figure 2 we "zoom-into" the black rectangle which spans the range from the lowest JWST masses currently observed up to the highest masses accessible via our numerical models. This is only possible for our outputs at $z=10.0$ and $z=15.0$ as halo masses large enough to be observed by JWST are simply not formed by $z=20.0$ (in a $\Lambda$CDM universe). Hence, we focus our analysis on the $z=15.0$ and $z=10.0$ outputs from here on out.
@@
The zoom-in region allows us to see what discrepancies do exist between the numerical HMFs and the (semi-)analytic fits at the masses and redshifts of most interest. At both $z=15.0$ and $z=10.0$ the numerical HMFs give a higher number density of haloes than any of the (semi-)analytic fits. In Figure 2 (top panel), we see a discrepancy between the Enzo and SWIFT results in the high-mass range at $z=15.0$. Previous work (e.g. Warren et al., 2006; More et al., 2011) has shown that the FOF mass is sensitive to mass resolution and the presence of substructure, which will be strongly dependent on redshift and which is difficult to correct for in general. It will also depend on local clustering, which can lead to distinct structures being linked by bridges of particles. This is the most likely explanation for the variance between Enzo and SWIFT at $z=15.0$. Note that this variance is absent from the HOP profiles (see Figure A.2).
In order to quantify any overall disagreements we make residual plots of all of the data in Figure 3 and Figure 4. In both figures we again show the entire mass range accessible at both redshift $z=15.0$ and $z=10.0$. In Figure 3 we show the variation between the Enzo simulations and the various (semi-)analytical fits. What we see is that the disagreement between the Enzo models and the fits varies between a factor of 2 (at both $z=10.0$ and $z=15.0$) and up to a factor of 100 (for the PS semi-analytic). Excluding the PS fits then, we see at most a difference of a factor of two between the Enzo (N-body) models and the fits. This is consistent with the results and conclusions of Yung et al. (2024).
In Figure 4 we show the equivalent plot from the SWIFT simulation i.e. we take the ratio of the HMF from the SWIFT outputs against the (semi-)analytic fits. At $z=10.0$ the agreement between the numerical (SWIFT) results and the analytic fits remains excellent. At $z=15.0$ we see a strong deviation towards higher masses which was not present for the Enzo simulations. It appears here that the Enzo runs are better able to resolve haloes at earlier times perhaps due to the inherent refinement strategy. A detailed analysis of the difference between the numerical codes is outside the scope of this paper and comparisons in this direction have been undertaken in the past (e.g. O’Shea et al., 2005; Regan, Haehnelt & Viel, 2007; Hayward et al., 2014).
@@
In summary the agreement between the numerical N-body solvers and the (semi-)analytic fits is excellent. In particular, the SMT fitting function gives excellent agreement with Enzo, with deviations of at most approximately a factor of two at $z=15.0$ (for low halo masses which are anyway currently unobservable) and converging to much less than a factor of two within the JWST window. For SWIFT , the agreement with SMT is equally excellent at $z=10.0$ but deviates somewhat for the higher halo masses at $z=15.0$. Using such (semi-)analytic fits then is unlikely to be a dominant source of error when testing high redshift observations against $\Lambda$CMD models.
@@ 4 Discussion
The recent explosion of data from the high-$z$ Universe, particularly beyond $z=10$, by JWST has led to a number of claims that the data is in tension with our galaxy formation and cosmological models (e.g. Boylan-Kolchin, 2023; Arrabal Haro et al., 2023; Yung et al., 2024; Finkelstein et al., 2024). The Universe beyond $z=10$ however is likely to be significantly different to the later and present day Universe. At $z\gtrsim 10$ galaxies are still in their infancy with the most massive galaxies at those epochs having stellar masses less than $10^{10}\ \rm{M}_{\odot}$ (and would be classified as dwarfs in the present day Universe). Moreover, there is strong evidence that the astrophysical processes at play at $z=10$ are sufficiently different to those of the present-day Universe and that they make significant alterations to the galaxy properties. This is particularly evident in galaxies like GN-z11 which is thought to harbour a massive black hole at it’s centre (Maiolino et al., 2024b), moreover this galaxy contains species abundances which are difficult to explain through standard processes (e.g. Bunker et al., 2023; Cameron et al., 2023; Charbonnel et al., 2023; Nandal et al., 2024). This peculiarity and lack-of-understanding is not unique to GN-z11 with a number of galaxies displaying properties which has evoked confusion within the community (e.g. Maiolino et al., 2024a).
@@
A key part of making progress in understanding the high-$z$ Universe is therefore to identify sources of systematic error in our models at high-$z$. In this paper we focus on exploring differences in fitting functions to the universal HMF and how they compare against direct N-body simulations at $z>10$. In particular we compare a wide range of fitting functions in use in the literature (see Table 4) against the adaptive mesh refinement code Enzo and against the N-body SPH code SWIFT.
@@
We find that for both Enzo and SWIFT that the match against the SMT analytic functions is excellent and within a factor of two up to $z=15.0$. The match against the standard Press & Schechter (1974) formalism is less accurate with deviations of up to an order of magnitude (at $z=10.0$ and $z=15.0$). Similarly, when comparing against standard fitting formula (see Table 4) we again see good agreement with deviations typically within an order of magnitude up to $z=15.0$ inside the window in which JWST can approximately observe high-z galaxies.
@@
We caution that the spatial resolution employed, controlled via the softening parameter for SWIFT and via the level of maximum refinement with Enzo , is set relatively high for our simulations. For example we use a softening length set to approximately the mean interparticle spacing divided by 25. This is slightly lower to the resolution we evolve the Enzo simulations with. This level of gravity resolution, specifically the softening length and spatial resolution here, is likely to be significantly higher than that used for typical galaxy formation simulations designed to run to $z\sim 0$ (e.g. Power et al., 2003, 2016; Zhang et al., 2019). It’s also worth noting that the gravitational collapse of dark matter halos is sensitive to the large scale gravitational field, which in a numerical simulation depends on the size of the simulation volume. Previous work (e.g. Power & Knebe, 2006) has highlighted how the mass function is sensitive to simulation volume, with a deficit of halos of a given mass at high masses. This effect is most pronounced in studies of the mass function at high redshifts, where the necessary high mass resolutions and large simulation volumes make these simulations particularly challenging. We therefore caution the reader that matching the results from numerical simulations designed primarily for large scale investigations may struggle to calculate the correct halo properties and abundances at $z\gtrsim 10$ (see e.g. Keller et al., 2023) unless the simulations are truly focused on high-$z$ study (see e.g. McCaffrey et al., 2023).
@@
Nonetheless, overall we find excellent agreement between N-body simulations and both analytic and fitting functions to HMFs, consistent with the results from Yung et al. (2024), and that these functions are unlikely to lead to large errors in our modelling of high-$z$ host haloes when appropriately modelled.
@@ Appendix A Alternative Halo Finding techniques
In addition to the widely used Friends-of-Friends Halo finding algoritm we also investigated how a different halo finder would impact our results. As discussed in §2 we also use the HOP halo (Eisenstein & Hut (1998)) in our analysis. Using the HOP halo finder we find near identical results to the FOF method and again we see approximately a factor of two difference between HOP and the analytic HMFs at both $z=10.0$ and $z=15.0$. The difference between HOP and the fitting functions varies up to an order of magnitude but similar to the FOF method the agreement is nonetheless still very good - particularly at $z=10.0$ where differences are typically less than a factor of two. In Figure A.1, Figure A.2, Figure A.3 and Figure A.4 we reproduce the figures from the main text but with the halo finding technique switched to HOP instead of FOF. We see negligible differences between FOF and HOP - a finding supported by other research studies (Knebe et al., 2011).
Figure A.1: Comparing the Enzo and SWIFT halo number densities contrasted with 5 halo number densities derived from fits at $z=20.0$ (left panel), $z=15.0$ (centre panel) and $z=10.0$ (right panel) using the HOP halo finder.
Figure A.2: High-mass sections of Figure A.1 (centre and right panels) shown in more detail.
Figure A.3: The ratios of the fitted halo number densities to the numerical halo number densities, derived from Enzo simulations and using the HOP halo finder.
Figure A.4: The ratios of the fitted halo number densities to the numerical halo number densities, derived from SWIFT simulations and using the HOP halo finder.

### v2

@@ 2.4.1 Press-Schechter Theory
where $\delta_{\text{c}}\approx 1.686$ is the critical overdensity required for a region to spherically collapse into a dark matter halo. A limitation of the PS fitting function is its tendency to overestimate the number of lower-mass halos and underestimate the number of higher-mass halos compared to N-body simulations (Lacey & Cole (1994), Sheth & Tormen (1999)). Since this seminal paper, many other fitting functions have been developed that aim to correct for this discrepancy as well as attempting to add additional sophistication to the modelling. In this work, we explore four other halo mass fitting functions in addition to PS.
@@ 2.4.2 Sheth, Mo & Tormen and Beyond
where $A=0.3222$, $a=0.707$ and $p=0.3$. We also test fitting functions developed by Reed et al. (2007) (Reed07) and Watson et al. (2013) (WatsonFoF refers to a fit using the Friends-of-Friends halo finder, WatsonSO refers to a fit using the Spherical Overdensity halo finder). The PS and SMT fits are widely-used and based on an analytical formalism, with no restriction on redshift or halo mass range. The Reed07, WatsonFoF and WatsonSO fits are based on simulation results, all with a redshift range of $z\,\in(0,30)$ and no explicit restriction on halo mass (see Table 4). While many other halo mass fitting functions exist, we choose these fits as they were all calibrated across the redshift range of interest to us (i.e. $z>10$). Other fitting functions are typically calibrated for lower redshifts.
@@ 3 Results
As previously stated, the goal of this paper is to compare halo number densities derived from simulation results with those derived from popular fits, and to test how well the fits can predict JWST halo mass abundances compared to simulations (e.g. Boylan-Kolchin, 2023).
Figure 4: The ratios of the fitted halo number densities to the numerical halo number densities, derived from Enzo simulations and using the FOF halo finder (with halo masses on a logarithmic scale). Differences between Enzo and the (semi-)analytical HMFs are typically less than a factor of two at $z=10.0$ increasing to a factor of 5 at $z=15.0$ (excluding the PS fit).
Figure 5: The ratios of the fitted halo number densities to the numerical halo number densities, derived from SWIFT simulations and using the FOF halo finder (with halo masses on a logarithmic scale). Similar to the Enzo result, SWIFT shows excellent agreement with the (semi-) analytic HMFs, all fits except PS agreeing within a factor of 2 for all but the highest masses at $z=10.0$. There is a greater discrepancy present at $z=15.0$, with some fits underestimating the numerical results by a factor of 100 at the high-mass range.
In Figure 2, we compare the HMFs from Enzo and SWIFT data (dashed lines) with the chosen fits (solid lines) for $z=20.0$, $z=15.0$ and $z=10.0$ using the FOF halo finder. The black rectangles indicate the high mass range depicted in more detail in Figure 3. The range of halo masses selected in the black rectangles is bounded by the approximate lowest halo masses observable by JWST at these epochs up to the most massive halo masses accessible by the numerical simulations at that redshift. This is only possible for our outputs at $z=10.0$ and $z=15.0$ as halo masses large enough to be observed (except perhaps via extreme lensing) by JWST are simply not formed by $z=20.0$ in a $\Lambda$CDM universe. Hence, we focus our analysis on the $z=15.0$ and $z=10.0$ outputs in particular. Overlaid onto Figures 2 to 5 are yellow rectangles which act as a visual aid, depicting an estimate for the range of halo masses that host recently-observed JWST galaxies and galaxy candidates. We choose sources with $15.010.0<z<$ 15.0  [^2] We use photometric redshift only if spectroscopic redshift is not yet available. and with an estimate of $M_{*}$ or $M_{\text{halo}}$ available. These sources include GN-z11 (Scholtz et al., 2024), Maisie’s Galaxy (Arrabal Haro et al., 2023b), GS-z14 (Helton et al., 2024) and others (Chakraborty et al., 2024). If there exists no estimate of $M_{\text{halo}}$ yet for a given source, we estimate it using the following equation:
@@
where we set the star formation efficiency parameter $\epsilon_{*}=0.1$. The width of the yellow rectangle is determined by the minimum and maximum halo masses in our sample and does not vary with redshift due to the small number of sources. In Figures 4 and 5, we take the ratios of the semi-analytic halo number densities against the Enzo and SWIFT outputs respectively to quantify the overall disagreements at $z=10.0$ and $z=15.0$.
@@
Over the full mass range, for both simulation suites and halo finders, we find the fits considered agree well across the entire spectrum. There is, over several dexes in mass and over the redshift range between $z=10.0$ and $z=20.0$, excellent agreement between all of the HMFs. Additionally, the fits improve as redshift decreases - not just between the numerical results and the (semi-)analytic results but even among the (semi-)analytic fits themselves.
At the lower mass range ($10^{6}$ to $10^{7}$ h-1 $\text{M}_{\odot}$), the semi-analytic fits overestimate the halo number densities by up to a factor of $\approx$ 2 compared to either Enzo or SWIFT data.
At the mid-range masses (107 to 109 h-1 M) at $z=10.0$, all but the PS fit overestimate the numerical simulations by less than a factor of 2. This is best illustrated by the bottom panels of Figures 4 and 5. For the same mass range at $z=15.0$, there is a greater discrepancy with all but the SMT and WatsonFOF fits underestimating the halo number densities by up to a factor of $\approx$ 5 compared to numerical simulations (see the top panels of Figures 4 and 5).
In the JWST mass range, we see the greatest discrepancies. At $z=15.0$, the PS semi-analytic fit underestimates the Enzo halo number density by a factor of $\approx$ 100 (see Figure 4 (top panel)). The other fits agree within a factor of $\approx$ 50 (WatsonSO) in the worst case with the SMT fit providing the closest match. We see a similar story at $z=10.0$ with a less extreme discrepancy (PS underestimating by a factor of $\approx$ 10, the other fits agreeing within a factor of $\approx$ 2 (see Figure 4 (bottom panel))). This is consistent with the results and conclusions of Yung et al. (2024).
@@
For the SWIFT simulations at $z=15.0$, we see a strong deviation towards higher masses which was not present for the Enzo simulations (see Figure 5 (top panel)). Both PS and WatsonSO underestimate the numerical halo number densities by a factor of $\approx$ 100, and SMT agrees best within a factor of $\approx$ 10. It appears here that the Enzo runs are better able to resolve halos at earlier times perhaps due to the inherent refinement strategy. A detailed analysis of the difference between the numerical codes is outside the scope of this paper and comparisons in this direction have been undertaken in the past (e.g. O’Shea et al., 2005; Regan, Haehnelt & Viel, 2007; Hayward et al., 2014). At $z=10.0$ the agreement between the numerical (SWIFT) results and the analytic fits remains excellent (see Figure 5 (bottom panel)). The PS fit underestimates the numerical results by a factor of $\approx$ 10, with the other fits agreeing within a factor of 2. In Figure 3 (top panel), we see a discrepancy between the Enzo and SWIFT results in the high-mass range at $z=15.0$. Previous work (e.g. Warren et al., 2006; More et al., 2011) has shown that the FOF mass is sensitive to mass resolution and the presence of substructure, which will be strongly dependent on redshift and which is difficult to correct for in general. It will also depend on local clustering, which can lead to distinct structures being linked by bridges of particles. This is the most likely explanation for the variance between Enzo and SWIFT at $z=15.0$. Note that this variance is absent from the HOP profiles (see Figure A.2).
@@
In summary the agreement between the numerical N-body solvers and the (semi-)analytic fits is excellent. In particular, the SMT fitting function gives excellent agreement with Enzo, with deviations of at most approximately a factor of 2 at $z=15.0$ (for low halo masses which are anyway currently unobservable) and converging to much less than a factor of two within the JWST window at $z=10.0$. For SWIFT , the agreement with SMT is equally excellent at $z=10.0$ but deviates somewhat for the higher halo masses at $z=15.0$. Using such (semi-)analytic fits, then, is unlikely to be a dominant source of error when testing high redshift observations against $\Lambda$CDM models.
@@ 4 Discussion
The recent explosion of data from the high-$z$ Universe, particularly beyond $z=10$, by JWST has led to a number of claims that the data is in tension with our galaxy formation and cosmological models (e.g. Boylan-Kolchin, 2023; Arrabal Haro et al., 2023a; Yung et al., 2024; Finkelstein et al., 2024). The Universe beyond $z=10$ however is likely to be significantly different to the later and present-day Universe. At $z\gtrsim 10$, galaxies are still in their infancy, with the most massive galaxies at those epochs having stellar masses less than $10^{10}\ \rm{M}_{\odot}$ (these would be classified as dwarfs in the present-day Universe). Moreover, there is strong evidence that the astrophysical processes at play at $z=10$ are sufficiently different to those of the present-day Universe and that they make significant alterations to the galaxy properties. This is particularly evident in galaxies like GN-z11 which is thought to harbour a massive black hole at its centre (Maiolino et al., 2024b), moreover this galaxy contains species abundances which are difficult to explain through standard processes (e.g. Bunker et al., 2023; Cameron et al., 2023; Charbonnel et al., 2023; Nandal et al., 2024). This peculiarity and lack-of-understanding is not unique to GN-z11 with a number of galaxies displaying properties which has evoked confusion within the community (e.g. Maiolino et al., 2024a). The most luminous galaxies observed by JWST remain in tension with state-of-the-art cosmological simulations (e.g. Keller et al. (2023)) with simulations struggling to model their extreme brightness at very early times. The reasons behind this are currently unknown but a greater emphasis on processes specific to the early Universe such as population III star formation and early black hole formation (e.g. McCaffrey et al. (2023)) may offer a pathway forward. It has also been argued that this tension may be resolved with the use of high-resolution simulations targeted at the high-redshift universe (see McCaffrey et al. (2023)).
@@
A key part of making progress in understanding the high-$z$ Universe is therefore to identify sources of systematic error in our models at high-$z$. Some sources of uncertainity in counting galaxies at such redshifts include cosmic variance (especially significant at the distances considered), error in stellar mass estimation and the presence of backsplash halos (halos that have lost some dark matter from the host halo, giving the impression of a higher baryon-dark matter ratio). These uncertainties have been explored thoroughly by Chen, Mo & Wang (2023). In this paper we focus on exploring differences in fitting functions to the universal HMF and how they compare against direct N-body simulations at $z\geq 10$. In particular, we compare a wide range of fitting functions in use in the literature (see Table 4) against the adaptive mesh refinement code Enzo and against the N-body SPH code SWIFT.
@@
We find that, for both Enzo and SWIFT , the match against the HMF (semi-)analytic functions is excellent, with many fits agreeing with numerical results within a factor of 2 for low and mid-range halo masses. The match against the standard Press & Schechter (1974) formalism is less accurate with deviations of up to an order of magnitude (at $z=10.0$ and $z=15.0$). Similarly, when comparing against standard fitting formula (see Table 4) we again see good agreement with deviations typically within an order of magnitude up to $z=15.0$ inside the window in which JWST can approximately observe high-$z$ galaxies.
@@
We caution that the spatial resolution employed, controlled via the softening parameter for SWIFT and via the level of maximum refinement with Enzo , is set relatively high for our simulations. For example, we use a softening length set to approximately the mean interparticle spacing divided by 25. This is slightly lower than the resolution we evolve the Enzo simulations with. This level of gravity resolution, specifically the softening length and spatial resolution here, is likely to be significantly higher than that used for typical galaxy formation simulations designed to run to $z\sim 0$ (e.g. Power et al., 2003, 2016; Zhang et al., 2019). It is also worth noting that the gravitational collapse of dark matter halos is sensitive to the large-scale gravitational field, which in a numerical simulation depends on the size of the simulation volume. Previous work (e.g. Power & Knebe, 2006) has highlighted how the mass function is sensitive to simulation volume, with a deficit of halos of a given mass at high masses. This effect is most pronounced in studies of the mass function at high redshifts, where the necessary high mass resolutions and large simulation volumes make these simulations particularly challenging. We therefore caution the reader that matching the results from numerical simulations designed primarily for large-scale investigations may struggle to calculate the correct halo properties and abundances at $z\gtrsim 10$ (see e.g. Keller et al., 2023) unless the simulations are truly focused on high-$z$ study (see e.g. McCaffrey et al., 2023).
@@
Nonetheless, overall we find excellent agreement between N-body simulations and both (semi-)analytic and fitting functions to HMFs, consistent with the results from Yung et al. (2024), and that these functions are unlikely to lead to large errors in our modelling of high-$z$ host halos when appropriately modelled.
Data availability
The analysis code used to produce the figures shown here is available upon request.
@@ Appendix A Alternative Halo Finding techniques
In addition to the widely-used Friends-of-Friends halo finding algorithm, we also investigate how a different halo finder would impact our results. As discussed in §2 we also use the HOP halo finder (Eisenstein & Hut (1998)) in our analysis. In Figures A.1 to A.4, we reproduce Figures 2 to 5 from the main text respectively but with the halo finding technique switched to HOP instead of FOF. Using the HOP halo finder we find near identical results to the FOF method and again we see approximately a factor of two difference between HOP and the analytic HMFs at both $z=10.0$ and $z=15.0$. The difference between HOP and the fitting functions varies up to an order of magnitude but similar to the FOF method the agreement is nonetheless still very good - particularly at $z=10.0$ where differences are typically less than a factor of two. We see negligible differences between FOF and HOP - a finding supported by other research studies (Knebe et al., 2011).
Figure A.1: Comparing the Enzo and SWIFT halo number densities contrasted with 5 halo number densities derived from fits at $z=20.0$ (upper panel), $z=15.0$ (centre panel) and $z=10.0$ (bottom panel) using the HOP halo finder (with halo masses and number densities on a logarithmic scale).
Figure A.2: High-mass sections of Figure A.1 (centre and bottom panels) shown in more detail.
Figure A.3: The ratios of the fitted halo number densities to the numerical halo number densities, derived from Enzo simulations and using the HOP halo finder (with halo masses on a logarithmic scale).
Figure A.4: The ratios of the fitted halo number densities to the numerical halo number densities, derived from SWIFT simulations and using the HOP halo finder (with halo masses on a logarithmic scale).